OpenStudy (anonymous):

Find an equation of variation in which Y varies as X and the following is true Y=40 and X=38

OpenStudy (gw2011):

y = kx y/x = k k is called the constant of variation Now putting in the values you get: 40/38 = k or 20/19 = k This is reduced to its lowest fraction from the above

OpenStudy (anonymous):

thanks but thats incorrect

OpenStudy (gw2011):

Sorry, but maybe I don't quite understand the question since I know that y=kx is an equation involving the constant of variation.